-3 and 5. are root of quadratic equation: find the equation
Found 3 solutions by josgarithmetic, laoman, richard1234:
Answer by josgarithmetic(39617)

(x+3)(x-5)=0

Answer by laoman(51)

To get the equation,
The general quadratic equation is
x^2 - (sum of roots)*x + product of roots
= x^2 - (-3+5)*x + (-3*5)
= x^2 - 2x -15

Answer by richard1234(7193)

k(x+3)(x-5) where k is a nonzero constant